# HGMF: A Hierarchical Gaussian Mixture Framework for Scalable Tool Invocation within the Model Context Protocol

### v1

@@ 4.1 Experiment Configuration
Fig. 3: Accuracy gain of the HGMF method over the MCP-zero baseline. The results show that HGMF’s performance advantage steadily increases with the sample size, demonstrating its superior scalability.
4.2 Result Analysis
@@
Table 1 and Fig. 2 present a comprehensive comparison of our proposed method, HGMF, against five baseline methods across eight different LLMs and four sample sizes. The results demonstrate that HGMF consistently achieves superior accuracy, particularly in high-shot settings. Fig. 3 shows the accuracy improvement of our proposed HGMF method over the MCP-zero baseline across eight LLMs and four sample sizes. Positive values indicate performance gains, while negative values indicate underperformance. Key observations are summarized as follows:
Overall Performance.
HGMF attains the highest average accuracy (82.34%) across all models, outperforming the strong baseline MCP-zero (78.10%) by over 4 percentage points. It achieves the best results on 6 out of 8 models, including Mistral:7b, Llama3:8b, Qwen:14b, Solar:10.7b, Openhermes:7b, and Phi3:3.8b. This confirms the effectiveness of our hierarchical clustering and pruning strategy in selecting high-quality tool candidates.
Comparison with Clustering-Based Baselines:
While K-means and Density-Based Clustering exhibit moderate improvements over Tokenization-Based sampling, they fall short of HGMF. Notably, Cluster-Weighted performs inconsistently, likely due to its reliance on local similarity scores without hierarchical context. In contrast, HGMF integrates both server-level and tool-level clustering, enabling more robust and context-aware selection.
Scalability with Sample Size:
HGMF’s performance gain increases consistently with more samples—from 41-shot (light gray) to 2797-shot (dark blue)—demonstrating strong scalability and its ability to extract high-quality candidates from larger tool pools.
Low-Shot Limitations:
In 41-shot settings, HGMF may underperform due to insufficient context for effective clustering. Notably, it lags behind MCP-zero on models like Gemma:7b and phi4:14b, with up to 30% accuracy drop.
High-Shot Superiority:
@@
At larger sample sizes ($\geq$401-shot), HGMF consistently outperforms the baseline. At 2797-shot, it delivers significant gains—over 40 percentage points on Openhermes:7b and 30+ points on phi4:14b, highlighting its strength in distilling relevant tools from noisy candidates.
Model Sensitivity:
Response to HGMF varies by model. Gemma:7b shows large fluctuation, while models like Mistral:7b and Qwen:14b exhibit stable improvements across all settings.
5 Conclusion
To address the challenges of context limitations and noise interference when Large Language Models select tools from large, hierarchical libraries, this paper introduces the Hierarchical Gaussian Mixture Framework (HGMF). The framework leverages the library’s inherent ”server-tool” structure, performing efficient probabilistic pruning on the vast toolset through two-stage Gaussian Mixture Model (GMM) clustering to produce a compact, high-relevance candidate set for the LLM.
Experiments confirm HGMF’s effectiveness and scalability. It achieves higher accuracy than baseline methods like random sampling across multiple LLMs. Critically, its performance advantage grows with the size of the tool library, demonstrating that the framework can effectively handle the challenges of large-scale, real-world toolsets. Through this intelligent filtering, HGMF not only improves selection accuracy but also reduces the LLM’s inference burden.
The core contribution of this work is an efficient and viable paradigm for large-scale tool invocation. By integrating hierarchical information into the pruning process, HGMF successfully decomposes a complex selection problem into more manageable sub-tasks. Although its performance is limited with very small toolsets, this work points toward future research, such as exploring adaptive clustering parameters and extending the framework to more complex, graph-structured tool libraries.

### v2

@@ 3.1. Experiment Configuration
Figure 4. Accuracy gain of the HGMF method over the MCP-zero baseline. The results show that HGMF’s performance advantage steadily increases with the sample size, demonstrating its superior scalability.
3.2. Result Analysis
Overall Performance
@@ 3.2. Result Analysis
Figure 2 present a comprehensive comparison of our proposed method, HGMF, against five baseline methods across eight different LLMs and four sample sizes. The results demonstrate that HGMF consistently achieves superior accuracy, particularly in high-shot settings. Figure 4 shows the accuracy improvement of our proposed HGMF method over the MCP-zero baseline. Positive values indicate performance gains, while negative values indicate underperformance. In summary, HGMF attains the highest average accuracy (82.34%) across all models, outperforming the strong baseline MCP-zero (78.10%) by over 4 percentage points. It achieves the best results on 6 out of 8 models. This confirms the effectiveness of our hierarchical clustering and pruning strategy in selecting high-quality tool candidates.
High-Shot Superiority
@@
As shown in Figure4, at larger sample sizes ,HGMF consistently outperforms the baseline. It delivers significant gains—over 40 percentage points on Openhermes:7b and 30+ points on phi4:14b. highlighting its strength in distilling relevant tools.
Effect of Regularization
An ablation study (Figure 3) confirms that our proposed regularization technique effectively mitigates performance instability on sparse data. While the baseline model without regularization struggles at low sample sizes (57-61% accuracy), the fully regularized HGMF yields an average accuracy gain of 21.65 percentage points in these low-shot regimes by preventing unstable clustering.
4. Conclusion
To tackle context limitations and noise in Large Language Models (LLMs) selecting tools from large, hierarchical libraries, this paper introduces the Hierarchical Gaussian Mixture Framework (HGMF). It exploits the library’s ”server-tool” structure, using two-stage Gaussian Mixture Model (GMM) clustering for efficient probabilistic pruning to generate a compact, highly relevant candidate set for the LLM. Experiments validate HGMF’s effectiveness and scalability, achieving higher accuracy than baselines like random sampling across various LLMs. Its advantage increases with tool library size, proving suitability for large-scale, real-world toolsets. By smart filtering, HGMF enhances selection accuracy and reduces LLM inference load. The core contribution is an efficient paradigm for large-scale tool invocation, integrating hierarchical information to break down complex selection into manageable sub-tasks. Though limited on very small toolsets, it suggests future directions like adaptive clustering parameters and extensions to graph-structured libraries.